ganeshie8 (ganeshie8):

power series of \(\sqrt{\dfrac{1-x}{1+x}}\)

OpenStudy (thomas5267):

Is there something I don't understand? Isn't power series just a Taylor series expansion? I don't think you need to ask this question with your mastery of math...

OpenStudy (thomas5267):

Something to do with spacetime diagrams?

ganeshie8 (ganeshie8):

not really, its just an approximation of doppler shift of light haha \[f = f_0\sqrt{\dfrac{1-\beta}{1+\beta}} = f_0(1-\beta+\frac{1}{2}\beta^2)\] assuming \(\beta \ll 1\)